OpenStudy (crashonce):

How do you find the maximum value of 2cos(A+10)cos(A+20)? What is the value of A which makes it a maximum

OpenStudy (crashonce):

@ganeshie8 @ikram002p sorry if ur eating lol

OpenStudy (ikram002p):

hehe yeah lol , xD im late to breakfast brb

ganeshie8 (ganeshie8):

can we write it as below ? \[\large 2\cos(A+10)\cos(A+20) = \cos(2A+30) + \cos(10)\]

OpenStudy (crashonce):

which identity?

ganeshie8 (ganeshie8):

2cosAcosB = cos(A+B) + cos(A-B)

OpenStudy (crashonce):

ohhhhhh ye ok

OpenStudy (anonymous):

a=30/2 ?

OpenStudy (crashonce):

i need steps @dg2 i dont care what the answer is, i need to learn

OpenStudy (anonymous):

idk i am asking....

ganeshie8 (ganeshie8):

rest should be easy to conclude as cos(10) is just a constant number

OpenStudy (crashonce):

ok but how do i find the max value

ganeshie8 (ganeshie8):

try below : \[\large -1\le \cos(2A+30) \le 1\]

ganeshie8 (ganeshie8):

\[\large -1+\cos(10)\le \cos(2A+30) + \cos(10)\le \color{purple}{1+\cos(10)}\]

ganeshie8 (ganeshie8):

that right side bound is the max value ^^

OpenStudy (anonymous):

a sin θ ± b cos θ = ±√ (a2 + b2 ) { for min. use – , for max. use + }

OpenStudy (anonymous):

when we are using this formulae idk?

OpenStudy (aum):

The question asks for the value of A that will make the expression a maximum.

OpenStudy (crashonce):

thanks guise

ganeshie8 (ganeshie8):

np :) again, the question is also asking about what value(s) of A makes it maximum...

ganeshie8 (ganeshie8):

you still need to solve : \( \cos(2A+30) = 1\) for A values that give u max

OpenStudy (crashonce):

ty

OpenStudy (anonymous):

this problem is over ahh?

OpenStudy (anonymous):

A plot and solution using Mathematica v9 is attached.